OpenStudy (hashuablue):

Can someone help me solve this system using elimination? 3x+2y=-17 x-3y=9

OpenStudy (anonymous):

Okay so what you want to do is combine the two equations into one long equation. So it looks like this: 3x+2y+x-3y= -17 + 9 From there, you want to combine like terms: \[3x+x+2y-3y = -17 + 9\] Finally, add/subtract the terms to reach your final answer: \[5x-y=8\] So \[5x-y=8\] is your final solution!